A toy rocket was launched from the ground. The function f(x) = -16x2 + 192x shows the height of the rocket f(x), in feet, from the ground at time x seconds.
What is the axis of symmetry of the graph of f(x), and what does it represent?
x = 12; it takes 12 seconds to reach the maximum height and 12 seconds to fall back to the ground
x = 12; it takes 12 seconds to reach the maximum height and 24 seconds to fall back to the ground
x = 6; it takes 6 seconds to reach the maximum height and 6 seconds to fall back to the ground
x = 6; it takes 6 seconds to reach the maximum height and 3 seconds to fall back to the groun

Answer by rothauserc(4718)

axis of symmetry is given by the following formula,
x = -b/2a
we are given,
x = -192 / (2*-16) = -192 / -32 = 6
x = 6; it takes 6 seconds to reach the maximum height and 6 seconds to fall back to the ground